integral of ( xdx)/(x^2+4)(x^2-5)^0.5

integral of (xdx)/(x^2+4)(x^2-5)^0.5

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Shivan ,

 ∫(-x dx) /(x2 + 4)(x2 -5)0.5  

substitute x2 -5 = t  we get xdx =tdt  substitute the two substitutions , 

∫-1/(t2+9) dt  = - tan-1( t/3) /3 + c

 resubstituting the substitution ,  I = - tan-1( (x2 -5) ^0.5 /3) /3 + c

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD